# Decentralized Stochastic Subgradient-based Methods for Nonsmooth Nonconvex Optimization

### v2

@@ 4 Application: Decentralized Stochastic Subgradient-based Methods
It is worth mentioning that Assumption 4.1 is mild and reasonable in practical scenarios. As demonstrated in [29, 28], the class of definable functions is wide enough to enclose a wide range of the loss functions of common nonsmooth neural networks. Moreover, [28] demonstrates that there exists a definable conservative field for any definable loss function constructed by definable blocks, within which the results yielded by Automatic Differentiation (AD) algorithms are contained.
Under Assumption 4.1 (1), we can choose the definable conservative field of $f_{i}$ as
$\mathcal{D}_{f_{i}}({\bm{x}}):=\frac{1}{|\mathcal{S}_{i}|}\sum_{l=1}^{|\mathcal{S}_{i}|}D_{F(\cdot;{\bm{s}}_{l})}({\bm{x}}).$ (4.1)
@@
Recall that $f({\bm{x}})=\frac{1}{d}\sum_{i=1}^{d}f_{i}({\bm{x}})$ as defined in (1.1), we can choose its definable conservative field $\mathcal{D}_{f}$ as
$\mathcal{D}_{f}({\bm{x}}):=\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\mathcal{D}_{f_{i}}({\bm{x}})\right).$ (4.2)
Then by [28, Theorem 5], we conclude that the set $\{{f({\bm{x}})}:{\bm{0}}\in D_{f}({\bm{x}})\}$ is a finite subset of $\mathbb{R}$.
@@ 4.1 Decentralized stochastic subgradient descent
6:         Compute ${\bm{d}}_{i,k}\in\frac{1}{|\mathcal{B}_{i,k}|}\sum_{l\in\mathcal{B}_{i,k}}D_{F(\cdot;{\bm{s}}_{l})}({\bm{x}})$;
@@
To establish the convergence properties of Algorithm 1, we first make the following assumptions based on Assumption 3.1.
Assumption 4.3.
@@
    There exists a constant $M_{x}>0$ such that
    $\sup_{k\geq 0}\left\|{\bm{X}}_{k}\right\|\leq M_{x}.$
@@
    The sequence of step-sizes $\{\eta_{k}\}$ satisfies
    $\lim_{k\to+\infty}\eta_{k}\log(k)=0,\text{ and }\sum_{i=0}^{+\infty}\eta_{k}=+\infty.$
Assumption 4.3(1) assumes the uniform boundedness of the sequence $\{{\bm{X}}_{k}\}=\{[\bm{x}_{1,k},...,\bm{x}_{d,k}]\}$. Moreover, Assumption 4.3(2) illustrates that the step-sizes diminish in the rate of $o(1/\log(k))$, hence enabling great flexibility in choosing the step-sizes $\{\eta_{k}\}$ in practical scenarios. In addition, as demonstrated in [30, Proposition 1.3], Assumption 4.3(2) implies Assumption 3.1(2) almost surely.
In the following theorem, we establish the global convergence of Algorithm 1 by showing that Algorithm 1 fits into our proposed framework with $\Phi=\mathcal{D}_{f}$, $\psi=f$ and $\mathcal{A}=\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$.
@@
Suppose Assumption 4.1 and 4.3 hold for Algorithm 1. Then for any $i\in[d]$, almost surely, any cluster point of $\{{\bm{X}}_{k}\}$ is a $\mathcal{D}_{f}$-critical point of (DOP) and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges.
@@
Let the filtration $\mathcal{F}_{k}=\sigma(\{{\bm{X}}_{l}:l\leq k\})$. Then we set ${\bm{Z}}_{k}={\bm{X}}_{k}$, ${\bm{H}}_{k}=\mathbb{E}[{\bm{D}}_{k}|\mathcal{F}_{k}]$ and $\Xi_{k+1}={\bm{D}}_{k}-\mathbb{E}[{\bm{D}}_{k}|\mathcal{F}_{k}]$. Then it is easy to verify that the update scheme (4.3) is equivalent to the update scheme (DSM).
Now we verify the validity of Assumption 3.1 and Assumption 3.3. Assumption 3.1(1) directly follows from Assumption 4.3(1) and local boundedness of the conservative field $D_{f_{i}}$ in (4.1). Moreover, Assumption 3.1(2) follows from Assumption 4.3(2) and [30, Proposition 1.3]. In addition, from the fact that $\mathbb{E}[{\bm{d}}_{i,k}|\mathcal{F}_{k}]\in\mathbb{E}_{\mathcal{B}_{i,k}}[D_{\mathcal{B}_{i,k}}({\bm{x}}_{i,k})]=D_{f_{i}}({\bm{x}}_{i,k})$, we can conclude that
@@
Therefore, by choosing $\Phi_{i}=\mathcal{D}_{f_{i}}$, we have verified the validity of Assumption 3.1(3).
Furthermore, according to [28, Section 6] and the path-differentiability of $f$, it follows that $f$ is the Lyapunov function of the differential inclusion $\frac{\mathrm{d}{\bm{x}}}{\mathrm{d}t}\in-\mathcal{D}_{f}({\bm{x}})$, and admits the stable set $\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$. In addition, Assumption 4.1(1) illustrates that $\{f({\bm{x}}):{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$ is finite in $\mathbb{R}$. As a result, we verify the validity of Assumption 3.3 with $\Phi=\mathcal{D}_{f}$, $\psi=f$, and $\mathcal{A}=\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in\mathcal{D}_{f}({\bm{x}})\}$.
Finally, from Theorem 3.9, we can conclude that almost surely, any cluster point of $\{{\bm{X}}_{k}\}$ is a $\mathcal{D}_{f}$-stationary point of (DOP) and the sequence of function values $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges. This completes the proof.
@@ 4.2 Decentralized generalized SGD with momentum
Momentum is a technique to accelerate the subgradient descent method in the relevant direction and dampen oscillations. In the single-agent setting, [36] investigates the global convergence properties of generalized SGD with momentum, where an auxiliary function $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is introduced to normalize the update directions of the primal variables. However, to the best of our knowledge, there is no existing work on developing its decentralized version.
In this subsection, motivated by the generalized SGD with momentum analyzed in [36], we first introduce the decentralized generalized SGD with momentum (DGSGDm) and present the details in Algorithm 2. Different from DSGD in Algorithm 1, Algorithm 2 introduces the auxiliary variables $\{{\bm{y}}_{i,k}:k\in\mathbb{N}\}$ to track the first-order moment of the stochastic subgradients of $f_{i}$. Therefore, Algorithm 2 requires two rounds of communication to perform local consensus to $\{{\bm{x}}_{1,k},\ldots,{\bm{x}}_{d,k}\}$ and $\{{\bm{y}}_{1,k},\ldots,{\bm{y}}_{d,k}\}$, respectively. Moreover, similar to the generalized SGD analyzed in [36], Algorithm 2 introduces an auxiliary function $\phi$ to normalize the update directions of its primal variables $\{{\bm{x}}_{1,k},\ldots,{\bm{x}}_{d,k}\}$.
Remark 4.5.
The auxiliary function $\phi$ in Algorithm 2 determines how the update directions of the variables $\{{\bm{x}}_{k}\}$ are regularized. Hence, Algorithm 2 yields different variants of DSGD-type methods with different choices of the auxiliary function $\phi$. For example, Algorithm 2 corresponds to DSGD-M when choosing $\phi({\bm{y}})=\frac{1}{2}\left\|{\bm{y}}\right\|^{2}$. Furthermore, when we select $\phi({\bm{y}})=\left\|{\bm{y}}\right\|_{1}$, Algorithm 2 employs the sign map to regularize the update directions of the variables $\{{\bm{x}}_{k}\}$. Under such selection of $\phi$, we can view this algorithm as a decentralized variant of SignSGD analyzed in [43], and denoted it as DSignSGD. Interested readers could refer to [36] for more details on the choices of the auxiliary function $\phi$.
Algorithm 2 DGSGDm for solving (DOP).
1:Initial point ${\bm{x}}_{0}=[{\bm{x}}_{1,0},\ldots,{\bm{x}}_{d,0}]$, initial descent direction ${\bm{y}}_{0}=[{\bm{y}}_{1,0},\ldots,{\bm{y}}_{d,0}]$, momentum parameter $\tau>0$, an auxiliary function $\phi:\mathbb{R}^{n}\to\mathbb{R}$, and a mixing matrix ${\bm{W}}$.
@@
8:         Compute ${\bm{d}}_{i,k+1}\in\frac{1}{|\mathcal{B}_{i,k+1}|}\sum_{l\in\mathcal{B}_{i,k+1}}D_{F(\cdot;{\bm{s}}_{l})}({\bm{x}})$;
@@
Let ${\bm{Y}}_{k}=[{\bm{y}}_{1,k},\ldots,{\bm{y}}_{d,k}]$ at $k$-th iteration in Algorithm 2, we present the following theorem to establish the global convergence of Algorithm 2.
@@
Suppose Assumption 4.1 and 4.3 hold for Algorithm 2. Moreover, we assume that the auxiliary function $\phi$ is convex and admits a unique minimizer at ${\bm{0}}$. Then for any $i\in[d]$, almost surely, any cluster point of $\{{\bm{X}}_{k}\}$ is a $\mathcal{D}_{f}$-stationary point of (DOP) and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges.
@@
Let the filtration $\mathcal{F}_{k}=\sigma(\{{\bm{X}}_{l},~{}{\bm{Y}}_{l}:l\leq k\})$. Then by setting ${\bm{Z}}_{k}=\left[\begin{smallmatrix}{\bm{X}}_{k}\\ {\bm{Y}}_{k}\end{smallmatrix}\right]$, ${\bm{H}}_{k}=\mathbb{E}\left[\left[\begin{smallmatrix}{\bm{T}}_{k}\\ \tau{\bm{Y}}_{k}{\bm{W}}-\tau{\bm{D}}_{k+1}\end{smallmatrix}\right]|\mathcal{F}_{k}\right]$, and $\Xi_{k+1}=\left[\begin{smallmatrix}{\bm{T}}_{k}\\ \tau{\bm{Y}}_{k}{\bm{W}}-\tau{\bm{D}}_{k+1}\end{smallmatrix}\right]-{\bm{H}}_{k}$, we can conclude that the update scheme (4.4) can be reshaped as ${\bm{Z}}_{k+1}={\bm{Z}}_{k}{\bm{W}}-\eta_{k}({\bm{H}}_{k}+\Xi_{k+1})$.
@@
Now we verify the validity of Assumption 3.1 and Assumption 3.3. From the uniform boundedness of $\{{\bm{X}}_{k}\}$ in Assumption 4.3(1) and the local boundedness of $\mathcal{D}_{f_{i}}$, we can conclude that $\{{\bm{D}}_{k}\}$ is uniformly bounded almost surely. Together with the update scheme of ${\bm{Y}}_{k}$, we can conclude that $\left\|{\bm{Y}}_{k+1}\right\|\leq(1-\tau\eta_{k})\left\|{\bm{Y}}_{k}\right\|+\tau\eta_{k}\left\|{\bm{D}}_{k+1}\right\|$, hence there exists a constant $M_{Y}>0$ such that $\sup_{k\geq 0}\left\|{\bm{Y}}_{k}\right\|\leq\max\{\left\|{\bm{Y}}_{0}\right\|,\sup_{k\geq 0}\left\|{\bm{D}}_{k}\right\|\}<M_{Y}$. Then combining with local boundedness of $\partial\phi$, it can be referred that $\{{\bm{T}}_{k}\}$ is uniformly bounded almost surely. This illustrates that Assumption 3.1(1) holds.
Moreover, by the definition of $\{\Xi_{k}\}$, it is straightforward to see that $\{\Xi_{k}\}$ is uniformly bounded almost surely. Then Assumption 4.3(2) and [30, Remark 1.5] illustrates that Assumption 3.1(2) holds almost surely for Algorithm 2. In addition, from the definition of $\mathcal{D}_{f_{i}}$, let $\delta_{k}=\left\|{\bm{X}}_{k+1}-{\bm{X}}_{k}\right\|_{\mathrm{F}}$ and $\Phi_{i}({\bm{x}},{\bm{y}})=\left[\begin{smallmatrix}\partial\phi({\bm{y}})\\ \tau{\bm{y}}-\tau\mathcal{D}_{f_{i}}({\bm{x}})\end{smallmatrix}\right]$, we can conclude that $\frac{1}{d}{\bm{H}}_{k}{\bm{1}}_{d}\in\frac{1}{d}\sum_{i=1}^{d}\Phi^{\delta_{k}}_{i}({\bm{x}}_{i,k},{\bm{y}}_{i,k})$. Followed by the same techniques in Lemma 3.5 and Proposition 3.6, we can prove that $\delta_{k}\to 0$. This verifies the validity of Assumption 3.1(3).

### v3

@@ 4 Convergence Properties of Decentralized Stochastic Subgradient-based Methods
As discussed in Section 2.3, the class of definable functions is wide enough to enclose the loss functions of common nonsmooth neural networks. Moreover, [36] demonstrates that there exists a definable conservative field for any definable loss function constructed by definable blocks, within which the results yielded by Automatic Differentiation (AD) algorithms are contained. Henceforth, Assumption 4.1(1) is mild and reasonable in practical scenarios.
For conciseness, we denote $D_{F_{i,l}}:=D_{F_{i}(\cdot;{\bm{s}}_{l})}$. According to [36, Corollary 4] and Assumption 4.1(1),
$D_{f_{i}}({\bm{x}}):=\frac{1}{|\mathcal{S}_{i}|}\sum_{l=1}^{|\mathcal{S}_{i}|}D_{F_{i,l}}({\bm{x}}).$ (4.1)
@@
is a definable conservative field for $f_{i}$. Recall that $f({\bm{x}})=\frac{1}{d}\sum_{i=1}^{d}f_{i}({\bm{x}})$ defined in (2.1), we can choose its definable conservative field $D_{f}$ as
$D_{f}({\bm{x}}):=\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}D_{f_{i}}({\bm{x}})\right).$ (4.2)
@@ 4.1 Decentralized stochastic subgradient descent
6:         Compute ${\bm{d}}_{i,k}\in\frac{1}{|\mathcal{B}_{i,k}|}\sum_{{\bm{s}}_{l}\in\mathcal{B}_{i,k}}D_{F_{i,l}}({\bm{x}}_{i,k})$;
@@
In Algorithm 1, the evaluation noise is introduced by the selection of mini-batch $\mathcal{B}_{i,k}$. Now, we discuss two different scenarios for selecting mini-batches and step-sizes.
Assumption 4.2 (Random reshuffling).
@@
    $\mathcal{B}_{i,k}$ is selected from set $\mathcal{S}_{i}$ by random reshuffling, meaning that for every epoch index $l\in\mathbb{N}$, the following holds:
    $\bigcup_{k=lN}^{(l+1)N-1}\mathcal{B}_{i,k}=\mathcal{S}_{i}.$
@@
    Let $\{\hat{\eta}_{k}\}$ be a prefixed sequence such that $\sum_{k=0}^{\infty}\hat{\eta}_{k}=+\infty,$ and $\lim_{k\to\infty}\hat{\eta}_{k}=0$. Moreover, for any $j_{1},j_{2}\in\mathbb{N}$, $\hat{\eta}_{j_{1}}=\hat{\eta}_{j_{2}}$ holds, if $\lfloor\frac{j_{1}}{N}\rfloor=\lfloor\frac{j_{2}}{N}\rfloor$. We set $\eta_{k}=c\hat{\eta}_{k}$, $c>0$ for all $k\in\mathbb{N}$.
Assumption 4.3 (With-replacement sampling).
  1. 1.
    $\mathcal{B}_{i,k}$ is selected from set $\mathcal{S}_{i}$ by with-replacement sampling for each $k\in\mathbb{N}$.
  2. 2.
    Let $\{\hat{\eta}_{k}\}$ be a prefixed sequence satisfying
    $\sum_{i=0}^{\infty}\hat{\eta}_{k}=+\infty,\text{ and }\lim_{k\to\infty}\hat{\eta}_{k}\log(k)=0.$
    We set $\eta_{k}=c\hat{\eta}_{k}$, $c>0$ for all $k\in\mathbb{N}$.
Assumption 4.2(1) and 4.3(1) employ different sampling techniques, leading to different forms of evaluation noise. In subsequent analysis, we will demonstrate that they conform to (3.8) and (3.18), respectively.
Assumption 4.2(2) and 4.3(2) both assume that $\{\eta_{k}\}$ is composed of a scaling parameter $c$ and a fixed diminishing sequence of step-sizes (at most in the rate of $o(1/\log(k))$), hence enabling great flexibility in choosing the step-sizes $\{\eta_{k}\}$ in practical scenarios.
In the following theorem, we establish the global convergence of Algorithm 1 by showing that Algorithm 1 fits into our proposed framework with $\Phi=D_{f}$, $\psi=f$ and $\mathcal{A}=\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in D_{f}({\bm{x}})\}$.
@@
Suppose Assumption 4.1 holds. Let $\{{\bm{X}}_{k}\}$ be the sequence generated by Algorithm 1.
  1. (1)
    When Assumption 4.2 holds, there exists $\alpha_{c}>0$, for any $c\in(0,\alpha_{c})$, it holds that almost surely, any cluster point of $\{{\bm{X}}_{k}\}$ is a $D_{f}$-critical point of (DOP) and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges for any $i\in[d]$.
  2. (2)
    When Assumption 4.3 holds, for any $\varepsilon>0$, there exists $\alpha_{c}>0$, for any $c\in(0,\alpha_{c})$,
    $\mathbb{P}\left(\lim_{k\to\infty}\mathrm{dist}({\bm{X}}_{k},\{{\bm{X}}\in\mathbb{R}^{n\times d}:{\bm{X}}={\bm{x}}{\bm{1}}^{\top},{\bm{0}}\in D_{f}({\bm{x}})\})=0\right)\geq 1-\varepsilon,$
    and
    $\mathbb{P}(f({\bm{x}}_{i,k})\textit{ converges })\geq 1-\varepsilon,\textit{ }\forall i\in[d].$
@@
Let the filtration $\mathcal{F}_{k}:=\sigma(\{{\bm{X}}_{l}:l\leq k\})$. We set ${\bm{Z}}_{k}:={\bm{X}}_{k}$, ${\bm{H}}_{k}:=\mathbb{E}[{\bm{D}}_{k}|\mathcal{F}_{k}]$ and $\Xi_{k+1}:={\bm{D}}_{k}-{\bm{H}}_{k}$. To demonstrate that Algorithm 1 fits into our framework (DSM), it suffices to check the update scheme (4.3) is a special case of (DSM) satisfying Assumption 3.1, and verify the validity of Assumption 3.8 or 3.12 with different sampling techniques. Assumption 3.1(1) follows immediately from Assumption 4.2(2) or 4.3(2); Assumption 3.1(2) follows quickly from local boundedness of the conservative field $D_{F_{i,l}}$. Nevertheless, how to verify (1.5) varies across different cases. For case (1) in Theorem 4.4, we set $\Phi_{i}:=D_{f_{i}}$. In this case, the nonnegative sequence $\{\delta_{k}\}$ in (1.5) is not subject to any other restrictions, hence we may select a sufficiently large one to meet the practical requirement for ${\bm{H}}_{k}$ and $\Phi_{i}$. This choice makes Assumption (1.5) hold trivially. For case (2), we need to carefully choose a diminishing $\{\delta_{k}\}$ so that it conforms to the conditions in Theorem 3.15.
Specifically, we note that $\mathbb{E}[{\bm{d}}_{i,k}|\mathcal{F}_{k}]\in\mathbb{E}_{\mathcal{B}_{i,k}}[\frac{1}{|\mathcal{B}_{i,k}|}\sum_{{\bm{s}}_{l}\in\mathcal{B}_{i,k}}D_{F_{i,l}}({\bm{x}}_{i,k})]=D_{f_{i}}({\bm{x}}_{i,k})$ in case (2), we can conclude that
@@
Therefore, by choosing $\Phi_{i}:=D_{f_{i}}$ and $\delta_{k}:=0$, we have verified the validity of (1.5).
Furthermore, according to [36, Section 6] and the path-differentiability of $f$, it follows that $f$ is the Lyapunov function of the differential inclusion $\frac{\mathrm{d}{\bm{x}}}{\mathrm{d}t}\in-D_{f}({\bm{x}})$, and admits the stable set $\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in D_{f}({\bm{x}})\}$. In addition, combining Assumption 4.1(1) and Proposition 2.7 yields $\{f({\bm{x}}):{\bm{0}}\in D_{f}({\bm{x}})\}$ is finite in $\mathbb{R}$. As a result, we verify the validity of Assumption 3.1 with $\Phi:=D_{f}$, $\psi:=f$, and $\mathcal{A}:=\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in D_{f}({\bm{x}})\}$.
[…]
The following theorem demonstrates that Algorithm 1 provably avoids spurious critical points. For clarity, we defer the full proof to Section A.
Theorem 4.5.
Suppose Assumption 4.1 holds. Let $\mathcal{X}_{0}$ be any compact set of $\mathbb{R}^{n}$, and sequence $\{{\bm{X}}_{k}\}$ be generated by Algorithm 1.
  1. (1)
    When Assumption 4.2 holds, there exists $\alpha_{c}>0$, a full-measure subset $\mathcal{S}$ of $(0,\alpha_{c})$, and a full-measure subset $\mathcal{X}_{1}$ of $\mathcal{X}_{0}$, such that for any $c\in\mathcal{S}$ and ${\bm{x}}_{0}\in\mathcal{X}_{1}$, it holds that almost surely, any cluster point of $\{{\bm{X}}_{k}\}$ is a Clarke-critical point of (DOP) and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges for any $i\in[d]$.
  2. (2)
    When Assumption 4.3 holds, for any $\varepsilon>0$, there exists $\alpha_{c}>0$, a full-measure subset $\mathcal{S}$ of $(0,\alpha_{c})$, and a full-measure subset $\mathcal{X}_{1}$ of $\mathcal{X}_{0}$, such that for any $c\in\mathcal{S}$ and ${\bm{x}}_{0}\in\mathcal{X}_{1}$,
    $\mathbb{P}\left(\lim_{k\to\infty}\mathrm{dist}({\bm{X}}_{k},\{{\bm{X}}\in\mathbb{R}^{n\times d}:{\bm{X}}={\bm{x}}{\bm{1}}^{\top},{\bm{0}}\in\partial f({\bm{x}})\})=0\right)\geq 1-\varepsilon,$
    and
    $\mathbb{P}(f({\bm{x}}_{i,k})\textit{ converges })\geq 1-\varepsilon,\textit{ }\forall i\in[d].$
@@ 4.2 Decentralized generalized SGD with momentum
Momentum is a technique to accelerate the subgradient descent method in the relevant direction and dampen oscillations. In the single-agent setting, [43] investigates the global convergence properties of generalized SGD with momentum, where an auxiliary function $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is introduced to normalize the update directions of the primal variables.
In this subsection, motivated by the generalized SGD with momentum analyzed in [43], we introduce the decentralized generalized stochastic subgradient descent method with momentum (DGSGD-M) and present the details in Algorithm 2. Different from DSGD in Algorithm 1, Algorithm 2 introduces the auxiliary variables $\{{\bm{y}}_{i,k}:k\in\mathbb{N}\}$ to track the first-order moment of the stochastic subgradients of $f_{i}$. Therefore, Algorithm 2 requires two rounds of communication to perform local consensus to $\{{\bm{x}}_{1,k},\ldots,{\bm{x}}_{d,k}\}$ and $\{{\bm{y}}_{1,k},\ldots,{\bm{y}}_{d,k}\}$, respectively. Moreover, similar to the generalized SGD analyzed in [43], Algorithm 2 introduces an auxiliary function $\phi$ to normalize the update directions of its primal variables $\{{\bm{x}}_{1,k},\ldots,{\bm{x}}_{d,k}\}$.
Algorithm 2 DGSGD-M for solving (DOP).
1:Initial point ${\bm{x}}_{0}\in\mathbb{R}^{n}$, initial descent direction ${\bm{y}}_{0}\in\mathbb{R}^{n}$, momentum parameter $\tau>0$, an auxiliary function $\phi:\mathbb{R}^{n}\to\mathbb{R}$, and a mixing matrix ${\bm{W}}$.
@@
8:         Compute ${\bm{d}}_{i,k+1}\in\frac{1}{|\mathcal{B}_{i,k+1}|}\sum_{{\bm{s}}_{l}\in\mathcal{B}_{i,k+1}}D_{F_{i,l}}({\bm{x}}_{i,k+1})$;
@@
Remark 4.6.
The auxiliary function $\phi$ in Algorithm 2 determines how the update directions of the variables $\{{\bm{x}}_{k}\}$ are regularized. Hence, Algorithm 2 yields different variants of DSGD-type methods with different choices of the auxiliary function $\phi$. For example, when we choose $\phi({\bm{y}})=\frac{1}{2}\left\|{\bm{y}}\right\|^{2}$, Algorithm 2 corresponds to DSGD with momentum; when we select $\phi({\bm{y}})=\left\|{\bm{y}}\right\|_{1}$, Algorithm 2 employs the sign map to regularize the update directions of the variables $\{{\bm{x}}_{k}\}$, which can be viewed as a decentralized variant of SignSGD analyzed in [51], and denoted as DSignSGD. Interested readers could refer to [43] for more details on the choices of the auxiliary function $\phi$.
Let ${\bm{Y}}_{k}:=[{\bm{y}}_{1,k},\ldots,{\bm{y}}_{d,k}]$ at $k$-th iteration in Algorithm 2. The convergence properties of this algorithm are established in Theorem 4.7, within which we concern about two different noise settings and step-sizes in Assumption 4.2 and 4.3.
@@
Suppose Assumption 4.1 holds. $\{{\bm{X}}_{k}\}$ is generated by Algorithm 2, and auxiliary function $\phi$ is convex and admits a unique minimizer at ${\bm{0}}$.
  1. (1)
    When Assumption 4.2 holds, there exists $\alpha_{c}>0$, for any $c\in(0,\alpha_{c})$, it holds that almost surely, any cluster point of $\{{\bm{X}}_{k}\}$ is a $D_{f}$-critical point of (DOP) and $\{f({\bm{x}}_{i,k}):k\in\mathbb{N}\}$ converges for any $i\in[d]$.
  2. (2)
    When Assumption 4.3 holds, for any $\varepsilon>0$, there exists $\alpha_{c}>0$, for any $c\in(0,\alpha_{c})$,
    $\mathbb{P}\left(\lim_{k\to\infty}\mathrm{dist}({\bm{X}}_{k},\{{\bm{X}}\in\mathbb{R}^{n\times d}:{\bm{X}}={\bm{x}}{\bm{1}}^{\top},{\bm{0}}\in D_{f}({\bm{x}})\})=0\right)\geq 1-\varepsilon$
    and
    $\mathbb{P}(f({\bm{x}}_{i,k})\textit{ converges })\geq 1-\varepsilon,\textit{ }\forall i\in[d].$
@@
Let the filtration $\mathcal{F}_{k}:=\sigma(\{{\bm{X}}_{l},~{}{\bm{Y}}_{l}:l\leq k\})$. Then by setting
${\bm{Z}}_{k}:=\left[\begin{smallmatrix}{\bm{X}}_{k}\\ {\bm{Y}}_{k}\end{smallmatrix}\right],{\bm{H}}_{k}:=\mathbb{E}\left[\left[\begin{smallmatrix}{\bm{T}}_{k}\\ \tau{\bm{Y}}_{k}{\bm{W}}-\tau{\bm{D}}_{k+1}\end{smallmatrix}\right]|\mathcal{F}_{k}\right],\text{ and }\Xi_{k+1}:=\left[\begin{smallmatrix}{\bm{T}}_{k}\\ \tau{\bm{Y}}_{k}{\bm{W}}-\tau{\bm{D}}_{k+1}\end{smallmatrix}\right]-{\bm{H}}_{k},$
we can conclude that the update scheme (4.4) can be reshaped as ${\bm{Z}}_{k+1}={\bm{Z}}_{k}{\bm{W}}-\eta_{k}({\bm{H}}_{k}+\Xi_{k+1})$, which aligns with the update format of (DSM).
@@
Now we check the validity of (1.5) and Assumption 3.1, and then show Assumption 3.8 and 3.12 are satisfied for case (1) and (2), respectively. From the local boundedness of $D_{f_{i}}$, we can conclude that $\{{\bm{D}}_{k}\}$ is uniformly bounded almost surely whenever $\{{\bm{X}}_{k}\}$ is bounded. Together with the update scheme of ${\bm{Y}}_{k}$, we can conclude that $\left\|{\bm{Y}}_{k+1}\right\|\leq(1-\tau\eta_{k})\left\|{\bm{Y}}_{k}\right\|+\tau\eta_{k}\left\|{\bm{D}}_{k+1}\right\|$, hence there exists a constant $M_{Y}>0$ such that $\sup_{k\geq 0}\left\|{\bm{Y}}_{k}\right\|\leq\max\{\left\|{\bm{Y}}_{0}\right\|,\sup_{k\geq 0}\left\|{\bm{D}}_{k}\right\|\}<M_{Y}$, if $\{{\bm{X}}_{k}\}$ is bounded. Then combining with local boundedness of $\partial\phi$, it can be referred that $\{{\bm{T}}_{k}\}$ is uniformly bounded almost surely whenever $\{{\bm{X}}_{k}\}$ is bounded. This illustrates that Assumption 3.1(2) holds.
Moreover, we set $\Phi_{i}({\bm{x}},{\bm{y}}):=\left[\begin{smallmatrix}\partial\phi({\bm{y}})\\ \tau{\bm{y}}-\tau D_{f_{i}}({\bm{x}})\end{smallmatrix}\right]$, then (1.5) holds automatically for case (1) with sufficiently large $\delta_{k}$. For case (2), we define $\delta_{k}$ as an exponential moving average of $\{\eta_{j}\|{\bm{T}}_{j}\|\}_{j\leq k}$, expressed as
$\delta_{k}:=\eta_{k}\|{\bm{T}}_{k}\|+2\sum_{j=0}^{k-1}\lambda_{2}^{k-1-j}\eta_{j}\|{\bm{T}}_{j}\|.$
From this, we can deduce that $\frac{1}{d}{\bm{H}}_{k}{\bm{1}}_{d}\in\frac{1}{d}\sum_{i=1}^{d}\Phi^{\delta_{k}}_{i}({\bm{x}}_{i,k},{\bm{y}}_{i,k})$ for case (2).